OpenStudy (anonymous):

write an equation in slope intercept form of the line that contains (-1,7) and (3,-9)..............................help me!!!!

OpenStudy (anonymous):

Hi!\[\Huge{\color{RED}{\mathbb{Welcome \quad To \quad OpenStudy!}}}\]

OpenStudy (anonymous):

Slope-intercept form: y = mx+b, where m is the slope of the line and b is the y-intercept. So use the equation above to find your answer just plug in the values given.

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

so will it be y=5/3x+-2

OpenStudy (anonymous):

Correct. replace that +- with just a -. y=5/3x-2

OpenStudy (anonymous):

thank you so much for your help and time!!!!

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

i just posted another question can you help me

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Post the link to your question on here. I can't find it on the list of questions. Is it in math?

OpenStudy (anonymous):

yes its http://openstudy.com/updates/50a68fe3e4b082f0b852b2be

OpenStudy (anonymous):

That's the link for this same question. If you need to post a new question, click "Close". Then you'll be able to ask another question.

OpenStudy (anonymous):

ok i did it it shoul be up there

OpenStudy (anonymous):

I'm sorry! Is this your question: write an equation in slope intercept form of the line that contains (-1,7) and (3,-9) ? I didn't notice you had edited your original question!

OpenStudy (anonymous):

First, find the slope. (-1,7) and (3,-9) \[m = \frac{ y2 - y1 }{ x2 - x }\]\[m = \frac{ -9 - 7 }{ 3 - (-1) } = \frac{ -9 - 7}{ 3+1 } = \frac{ -16 }{ 4 }= -4\]Slope is -4 -------------------- You have a point (-1, 7) and the slope(-4) so use the Point-Slope formula. y - y1 = m(x - x1) y - 7 = -4(x - (-1)) y - 7 = -4(x + 1) \[Distrbute\]y - 7 = -4x - 4 \[Add \quad 7\]y = -4x + 3